OpenStudy (anonymous):

Simplify each expression. −− √63t^4

OpenStudy (anonymous):

3t^2sqrt7

OpenStudy (anonymous):

\[\sqrt{63 \times t ^{4}} = \sqrt{63} \times \sqrt{t^{4}} = \sqrt{9 \times 7} \times \sqrt{t^{4}} =...\] \[3 \sqrt{7}\ t^{2}\]